### v4

# Certificate-Sensitive Subset Sum: Realizing Instance Complexity

Abstract
The Subset Sum problem is a classical NP-complete problem with a long-standing $O^{*}(2^{n/2})$ deterministic bound due to Horowitz and Sahni. Yet, many instances are much easier, as their complexity is governed by $U=|\Sigma(S)|$, the number of unique subset sums. We present three contributions that exploit this structure. First, we introduce the first deterministic algorithm whose runtime provably grows with the certificate size $U$, constructing all distinct subset sums in time $O(U\cdot n^{2})$, with a randomized variant achieving expected time $O(U\cdot n)$. This provides a constructive link to Instance Complexity. Second, we build a solver that combines this enumerator with a double meet-in-the-middle strategy and a Controlled Aliasing technique, achieving a worst-case runtime of $O^{*}(2^{n/2-\epsilon})$ for a constant $\epsilon>0$—the first deterministic improvement over the classical bound on all instances. Finally, we refine fine-grained hardness for Subset Sum by showing that SETH-based reductions implicitly assume collision-free instances with maximal $U$.
1 Motivation
Our work delivers a structural breakthrough on three fronts.
  • A first in certificate-sensitive complexity: This is, to our knowledge, the first deterministic algorithm for a canonical NP–complete problem whose runtime adapts to a provably minimal certificate, namely ($\Sigma(S)$). Our algorithm constructs this certificate online from scratch, rather than merely deciding feasibility.
  • Strictly faster than Horowitz–Sahni on all instances: IC-SubsetSum (formerly Beyond Worst-Case Subset Sum from [11]) incorporates a novel pruning rule that guarantees worst-case runtime $O^{*}(2^{n/2-\varepsilon})$ for a constant $\varepsilon>0$. It is the first deterministic solver to provably outperform the classical $O^{*}(2^{n/2})$ bound on every input, without exception.
  • A structural reframing of fine-grained lower bounds: Existing fine-grained reductions rely on the assumption that hard instances have $U=2^{n/2}$ (i.e., no subset sums collide). IC-SubsetSum shows that this assumption is not just technical, but structural: the runtime hardness of Subset–Sum is a direct function of its subset sum collisions.
  • A generic approach for certificate-sensitive algorithms: By reinterpreting dynamic programming as a guided certificate traversal, IC-SubsetSum offers a new design template for adaptive enumeration. This opens the door to structural algorithms for other NP–complete problems where certificate size varies widely across instances.
2 Introduction
Worst–Case versus Per–Instance Analysis.
The modern theory of algorithms is dominated by worst–case running–time guarantees. While immensely successful, the paradigm sometimes fails to explain the vast performance variance that practitioners observe between individual inputs of the same size. A complementary research line, initiated by [10], formalised Instance Complexity (IC): the intrinsic difficulty of a single instance is the bit–length of the shortest “special–case program” that decides it within a time bound $t$.
At first glance, however, IC seems paradoxical: for each among infinitely many possible inputs we would have to synthesise a bespoke decision program before seeing that input, with no a priori structural knowledge. Lacking such a generative recipe, researchers long treated IC as a purely existential benchmark—useful for lower bounds but incompatible with efficient algorithms. IC-SubsetSum challenges this perception by constructing the certificate online and bounding its runtime by a provable function of the certificate length.
The Subset–Sum Problem.
In the canonical Subset–Sum problem we are given a multiset $S=\{a_{1},\ldots,a_{n}\}\subseteq\mathbb{Z}_{>0}$ and a target $t\in\mathbb{Z}_{>0}$. The task is to decide whether some submultiset sums to $t$. Subset–Sum is NP–complete and underpins numerous reductions across combinatorics, cryptography, and fine–grained complexity.
Classical Algorithms.
Two textbook techniques illustrate the gulf between worst–case and per–instance behaviour:
  1. a.
    Bellman Dynamic Programming takes $O(nt)$ time—pseudopolynomially efficient when $t$ is small, but infeasible for large numeric ranges [3]. While a small target $t$ naturally constrains $U$, the reverse is not true; our approach remains efficient even for large $t$ provided that $U$ is small due to additive structure.
  2. b.
    Horowitz–Sahni Meet–in–the–Middle enumerates all $2^{n/2}$ subset sums of each half of $S$ and intersects the two lists in $O^{*}(2^{n/2})$ time [8]. Decades of work have failed to beat the $2^{n/2}$ factor in the worst case [4, 16].
Yet practitioners notice that Subset–Sum instances stemming from real data (e.g. knapsack cryptosystems, bin–packing logs) are routinely solvable in far less time. The explanation is simple: such instances exhibit enormous collision rate—many distinct subsets happen to yield identical sums. Classical analyses, parameterized solely by $n$ and $t$, overlook this structure.
Bridging Theory and Practice with IC-SubsetSum.
[…]
This work builds upon the algorithm initially introduced in the ‘Beyond Worst-Case Subset Sum’ framework [11], now formalized under the IC-SubsetSum model.
Contributions.
Our work makes four contributions:
  • Certificate–sensitive enumeration. We design and analyse IC-SubsetSum, the first deterministic algorithm to construct the certificate $\Sigma(S)$ with a runtime of $O(U\cdot n^{2})$ that adapts to the instance’s structure. We also present a randomized variant that achieves an expected runtime of $O(U\cdot n)$.
  • A guaranteed worst–case improvement. We prove that IC-SubsetSum runs in $O^{*}(2^{n/2-\varepsilon})$ time for a fixed constant $\varepsilon>0$, making it strictly faster than classical meet-in-the-middle algorithms on all inputs, regardless of structure.
  • Refined lower–bound methodology. We formalise a “collision–free” promise variant of Subset–Sum and show that classical ETH/SETH reductions only rule out $2^{n/2}$ algorithms on that structure–resistant slice. Future reductions must certify low collision entropy.
  • Template for other problems. We discuss how the certificate–sensitive viewpoint can guide new algorithms for knapsack, partition, and beyond.
Advantage over Dynamic Programming.
The certificate-sensitive bound provides an exponential advantage over classical dynamic programming (DP) on instances with large numeric range but low structural complexity. For example, consider an instance $S$ with $n-1$ elements equal to 111, and one large element $L=2^{n}$. For a target $t=L$, the DP runtime is $O(n\cdot t)=O(n\cdot 2^{n})$. In contrast, the number of unique subset sums for $S$ is only $U=O(n)$, yielding a polynomial runtime of $O(U\cdot n^{2})=O(n^{3})$ for IC-SubsetSum. Furthermore, the underlying enumeration framework supports anytime and online operation, allowing for interruption and incremental updates; these features are detailed in our companion technical paper [11].
Relation to Companion Papers.
This paper lays the foundational theoretical groundwork for the Certificate-Sensitive framework, building upon the algorithmic and empirical results presented in our technical companion paper [11]. The framework’s approach is shown to be broadly applicable in two other companion papers, released concurrently, that extend it to the 0–1 Knapsack problem [13, forthcoming] and the 3-SAT problem [12, forthcoming].
Organisation.
Section 3 introduces preliminaries. Section 4 presents our first contribution: a certificate-sensitive framework that algorithmically realizes Instance Complexity. Section 5 develops our second contribution, a deterministic Subset Sum solver with a worst-case runtime below the classical $2^{n/2}$ bound. Section 6 contributes a refined structural perspective on fine-grained hardness and its implicit assumptions. Section 7 provides empirical validation. Finally, Sections 810 place our results in broader context and outline directions for future research.
3 Preliminaries
Notation.
Let $S=\{a_{1},\ldots,a_{n}\}\subseteq\mathbb{Z}_{>0}$ be a multiset of $n$ positive integers. We write $[n]=\{1,2,\ldots,n\}$ and let
$\Sigma(S)=\left\{\sum_{i\in T}a_{i}\;\middle|\;T\subseteq[n]\right\}$
denote the set of distinct subset sums. The empty sum is included by convention, so $0\in\Sigma(S)$. Let $U:=|\Sigma(S)|$. Throughout this paper, we consider each sum $\sigma\in\Sigma(S)$ to be implicitly paired with its canonical prefix representation (i.e., its lexicographically minimal index mask).
Subset Sum Notation.
For any subset $A\subseteq S$, we write $\sigma(A)$ to denote the sum of its elements:
$\sigma(A):=\sum_{a\in A}a.$
This distinguishes numerical subset sums from the set of all distinct sums, denoted $\Sigma(S)$.
Definition 1 (Distinct Subset Sums).
The quantity $U=|\Sigma(S)|$ denotes the number of distinct subset sums of $S$. Trivially, $1\leq U\leq 2^{n}$. The lower bound is met when $S$ is empty, and the upper bound is met when all $2^{n}$ subset sums are unique.
This parameter $U$ will serve as our central structural complexity measure. It governs both the algorithmic behavior of IC-SubsetSum and the size of the certificate it produces.
Collision Entropy.
We define the (base-2) collision entropy of $S$ as
$H_{c}(S):=n-\log_{2}|\Sigma(S)|=n-\log_{2}U.$
This quantity measures the compressibility of the subset sum space. Intuitively, $H_{c}(S)$ captures how much smaller $\Sigma(S)$ is than the full power set: when many subsets collide onto the same sum, $U\ll 2^{n}$ and $H_{c}(S)$ is large; when most sums are unique, $U\approx 2^{n}$ and $H_{c}(S)$ is near zero.
  • If all $a_{i}=1$, then every subset sum lies in $\{0,1,\ldots,n\}$ and $U=n+1=O(n)$, so $H_{c}(S)=n-\log_{2}n=\Theta(n)$.
  • If $S$ is constructed to be collision-free (e.g., a superincreasing sequence), then $U=2^{n}$, so $H_{c}(S)=0$ and no structure is exploitable; IC-SubsetSum runs in worst-case time.
Collision entropy may be viewed as a coarse proxy for the Kolmogorov compressibility of the subset sum landscape: the more structured the set, the lower its information content.
Role of $U$ in Complexity.
The parameter $U=|\Sigma(S)|$ governs both the structural difficulty of a Subset Sum instance and the size of its minimal certificate. If $\Sigma(S)$ is known, then deciding whether some $T\subseteq S$ sums to a target $t$ reduces to checking whether $t\in\Sigma(S)$—a constant-time membership query. Thus, $U$ captures the certificate length for the decision variant.
For the constructive variant, a solution subset must be recovered; here, each subset sum must be paired with a canonical encoding of the realizing subset (e.g., a bitmask). Our algorithm, IC-SubsetSum, builds this richer certificate explicitly, and its runtime and space scale with $U$ and $n$. Designing algorithms whose performance adapts to $U$, rather than worst-case size $2^{n}$, is the central goal of this work.
Certificates for Subset Sum.
Once the set $\Sigma(S)$ is known, deciding whether there exists a subset of $S$ summing to a target $t$ reduces to checking if $t\in\Sigma(S)$. Thus, $\Sigma(S)$ serves as an information-theoretically minimal certificate for the decision version of Subset Sum: no smaller object suffices to resolve all yes/no queries about subset-sum feasibility.
However, for the constructive version—retrieving an actual subset $T\subseteq S$ such that $\sum_{i\in T}a_{i}=t$—this is no longer sufficient. In this case, each sum must be paired with a compact encoding of a corresponding witness subset. Our algorithm constructs such a certificate by maintaining, for each $\sigma\in\Sigma(S)$, a canonical bitmask representing the lex-minimal subset that realizes $\sigma$.
  • The decision certificate is $\Sigma(S)$, with total length $O(U\cdot\log(n\cdot\max(S)))$ bits if each sum is represented in binary.
  • The constructive certificate is the set of $(\sigma,\text{bitmask})$ pairs, with total size $O(U\cdot n)$ bits.
This distinction underlies the runtime and space guarantees of our algorithm, which produces the stronger constructive certificate online. All references to “certificate size” henceforth will clarify which variant is being discussed.
Instance Complexity (IC).
Instance Complexity [10] formalises the idea that hard problems can have easy instances—even without randomization or approximation. The IC of an instance $x$ under time bound $t$ is defined as the bit-length of the shortest program $P$ that outputs the correct answer on $x$ within time $t$. It is denoted $\mathsf{IC}_{t}(x)$.
For Subset–Sum, the list $\Sigma(S)$ plays a dual role. It is not only a certificate but also a low-complexity proxy for a correct decision procedure. Once $\Sigma(S)$ is known, any target $t$ can be resolved in time $O(1)$ via a membership query. Hence, $\Sigma(S)$ encodes a special-case program of length $O(U\log(n\cdot\max(S)))$, yielding a concrete upper bound on $\mathsf{IC}_{t}((S,t))$. When the goal is to construct a solution, the richer certificate consisting of $(\sigma,\text{bitmask})$ pairs provides a special-case program of length $O(U\cdot n)$.
In our IC discussion, we assume a time bound $t(n)=O(1)$ when $\Sigma(S)$ is stored in a hash table, or $t(n)=O(\log U)$ if stored in sorted order for binary search. Both bounds reflect the efficient nature of checking membership in a precomputed certificate set.
4 Certificate-Sensitive Framework
4.1 Unique-Subset-Sum Enumerator
Our algorithm begins by generating $\Sigma(S)$, the set of all distinct subset sums of $S$. The core challenge is to do this without duplication—i.e., without computing the same sum multiple times via different subsets. We use an ordered traversal strategy based on prefix extensions.
We emphasize that the goal of the enumerator is not merely to determine the feasibility of subset sums, but to generate the full constructive certificate as defined in Section 3. That is, for each unique subset sum $\sigma\in\Sigma(S)$, the algorithm records the lexicographically minimal bitmask of a subset realizing $\sigma$. This allows efficient reconstruction of actual witnesses and is necessary for constructive queries. Consequently, the runtime and space complexity of the enumerator will scale with the size of this richer certificate. Full implementation details and pseudocode for the enumerator are provided in Appendix A.
Canonical Prefixes for High-Density Instances.
In high-density instances, where many distinct subsets may sum to the same value, a naive ”first-prefix-wins” pruning strategy can fail. To guarantee correctness, our algorithm resolves collisions deterministically by defining a canonical prefix for each sum—this is the prefix whose index mask is lexicographically minimal among all subsets producing that sum. By retaining only this canonical representation when multiple extensions lead to the same sum, we suppress redundant work and ensure each unique sum is discovered via a single, well-defined path.
Bitmask Representation.
The algorithm represents each prefix by a bitmask of length $n$, where the $i$-th bit indicates whether $a_{i}$ is included. Bitmask comparison to determine the lex-minimum takes $O(n)$ time. The total number of unique sums is $U$, so we store at most $U$ bitmasks.
Theorem 1 (Deterministic Runtime of the Enumerator).
The deterministic enumerator computes $\Sigma(S)$ in $O(|\Sigma(S)|\cdot n^{2})$ worst-case time and requires $O(|\Sigma(S)|\cdot n)$ space.
Proof.
The runtime is dominated by the prefix extension and deduplication process. There are $U=|\Sigma(S)|$ unique sums, or ”states,” to be discovered. From each state, the algorithm attempts to extend its prefix with one of at most $n$ elements, resulting in $O(U\cdot n)$ extension attempts. In the event of a sum collision, the algorithm must select a canonical prefix by comparing two $n$-bitmasks, which takes $O(n)$ time. The total runtime is therefore $O(U\cdot n)\times O(n)=O(U\cdot n^{2})$. The space complexity is determined by storing the $n$-bitmask for each of the $U$ states, resulting in $O(U\cdot n)$ space usage. ∎
Theorem 2 (Randomized Runtime of the Enumerator).
A randomized variant of the enumerator computes $\Sigma(S)$ in expected time $O(|\Sigma(S)|\cdot n)$.
Proof.
The randomized variant replaces the $O(n)$ prefix comparison with an expected $O(1)$ comparison using polynomial hashing (e.g., from a 2-universal hashing family) with a randomly chosen seed. This ensures that, with high probability, no two distinct bitmasks collide. Thus, the expected comparison cost per extension remains $O(1)$ even under adversarial inputs. ∎
Remark 1.
The expected runtime bound for the randomized variant assumes a standard cryptographic model in which the hash function is drawn from a universal family. This ensures robustness against adversarially chosen inputs by keeping the probability of bitmask collisions negligible. To ensure a negligible collision probability (e.g., $1/\mathrm{poly}(n,U)$), the polynomial hash function is evaluated over a prime field of size $\mathrm{poly}(n,U)$.
4.2 Discussion and Conditional Optimality
A Local Enumeration Model.
We define a natural model of computation that captures the core constraints of real-time certificate construction. The local deduplicating enumeration model assumes that the algorithm:
  • maintains a growing set $P$ of seen prefixes, each encoding a valid subset sum and its associated path;
  • in each round, selects a candidate prefix $p\in P$ and a next element $a_{i}$ to extend it with;
  • uses local information about $p$ and $a_{i}$ to decide whether to emit the new sum or prune it.
In this setting, the main computational bottleneck arises in testing whether two prefixes yield the same sum and comparing their canonicality.
A Conditional Lower Bound.
We now formalize a mild but powerful conjecture.
Conjecture 1.
In the local deduplicating enumeration model, any algorithm that emits all $U$ distinct subset sums must take $\Omega(U\cdot n)$ time in the worst case.
This conjecture posits that the $O(U\cdot n)$ benchmark is the best we can hope for in general, even when allowing randomization. It reflects the intuition that each solution must be “seen” at $n$ bits of resolution to decide whether to include it.
Theorem 3 (Conditional Optimality of Randomized IC-SubsetSum).
If Conjecture 1 holds, then the expected $O(U\cdot n)$ runtime of the randomized IC-SubsetSum algorithm is optimal in expectation within the local deduplicating model.
Conclusion.
We believe this represents the first instance-sensitive enumeration algorithm for Subset Sum whose performance is provably tied to the certificate size $U$. Our randomized variant matches the conjectured optimal runtime of $O(U\cdot n)$ in expectation. A key open question remains whether a deterministic algorithm can also achieve this bound, or whether the additional $O(n)$ overhead for comparison—leading to an $O(U\cdot n^{2})$ runtime—is inherent for any deterministic approach in this model.
4.3 Constructive Link to Instance Complexity
Instance complexity, introduced by Orponen, Ker-i, Uwe and Watanabe [10], measures the minimum information needed to decide a single input instance $x$ of a decision problem . Let $\mathsf{P}$ be the class of polynomial-time decidable languages.
Definition 2 (IC, informal).
The instance complexity of $x\in\{0,1\}^{n}$ with respect to $L$ is the bit-length of the shortest program $P_{x}$ such that:
  • $P_{x}$ halts on every input $y$ of size $n$ within time $t(n)$,
  • $P_{x}(x)=L(x)$, and
  • $P_{x}(y)$ may be undefined or incorrect when $y\neq x$.
They showed that for some languages in NP, the IC of a random input can be exponentially smaller than any universal algorithm. Yet because IC is defined existentially, not constructively, it was long thought to offer no algorithmic advantage.
Proposition 1 (from [10]).
Let $L\in\mathsf{NP}$. Then for every polynomial-time verifier $V$ for $L$, the IC of a yes-instance $x\in L$ is at most the bit-length of its shortest witness.
IC-SubsetSum realizes instance complexity in a relaxed model where the constructed program is permitted to be undefined or incorrect on non-target inputs. To comply with the canonical definition of Instance Complexity [10], which requires partial correctness (i.e., the program must return either the correct answer or $\bot$ on all inputs), the certificate $\Sigma(S)$ must be embedded within a wrapper that first checks whether the input set matches the specific set $S$ for which the certificate was generated. If the check fails, the program returns $\bot$; otherwise, it uses $\Sigma(S)$ to decide membership for the provided target $t$. This adjustment preserves correctness but incurs an additive overhead of $O(|S|)$ in program size. Thus, while the runtime remains governed by $U=|\Sigma(S)|$, the size of the canonical IC program is $O(U\cdot n+|S|)$.
IC-SUBSETSUM as a Certifier.
As discussed above, the raw certificate $\Sigma(S)$ must be paired with an input-checking wrapper to form a canonical IC program. What distinguishes IC-SubsetSum is that it does not merely verify membership in $\Sigma(S)$; it constructs the entire certificate from scratch. Our deterministic algorithm achieves this in $O(U\cdot n^{2})$ worst-case time, while our randomized variant does so in $O(U\cdot n)$ expected time.
Implication
[…]
While Instance Complexity is classically defined in a non-uniform setting—allowing a different short program for each instance—IC-SubsetSum provides a uniform, deterministic algorithm that adapts to the structure of each input. It synthesizes the decision procedure on the fly, realizing the promise of IC in an algorithmically constructive form.
This resolves the long-standing tension between the theoretical power of instance complexity and its historically non-constructive formulation.
[…]
5.1 Double Meet-in-the-Middle Solver
We now describe how to solve Subset–Sum given only the enumerators for $\Sigma(S)$. The detailed pseudocode for the full solver is presented in Appendix C.
Clarifying Certificate Scope.
While the unique-subset-sum enumerator described above can be applied to the full input $S$ to construct the complete certificate $\Sigma(S)$ in time $O(U\cdot n^{2})$, our solver employs a more efficient strategy. It splits $S$ into halves $\ell_{0}$ and $\ell_{1}$ and applies the enumerator separately to each side, yielding certificates $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$ with total runtime $O(\max(U_{0},U_{1})\cdot n^{2})$. This avoids ever constructing $\Sigma(S)$ explicitly. The solver answers the query $t\in\Sigma(S)$ by testing cross-split combinations on-the-fly. It trades away post hoc queryability: to check a new target $t^{\prime}$, the merge logic must be repeated. Throughout the paper, we use $U=|\Sigma(S)|$ as a global proxy for instance complexity, but emphasize that our solver’s runtime depends operationally on $U_{0}$ and $U_{1}$.
Splitting the instance.
Let $S=\ell_{0}\cup\ell_{1}$ be a partition into left and right halves. We enumerate $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$ using the prefix-unique method described above. Let $U_{0}=|\Sigma(\ell_{0})|$ and $U_{1}=|\Sigma(\ell_{1})|$. Note that the additive structure within each half can lead to a significant imbalance (e.g., $U_{0}\ll U_{1}$), though this does not affect the overall asymptotic bound.
Solving and certifying.
To decide whether any subset of $S$ sums to $t$, we check for each sum $x\in\Sigma(\ell_{0})$ whether a corresponding match can be found in $\Sigma(\ell_{1})$. This check is more comprehensive than a simple search for $t-x$, as it also considers complements and mixed cases to cover all solution structures, as detailed in the lemmas of Appendix B. To certify all solutions, we can track all such valid $(x,y)$ pairs and output their associated bitmasks.
Theorem 4 (Certificate-Sensitive Solver).
There exists a deterministic algorithm that solves Subset–Sum in time $O(\max(U_{0},U_{1})\cdot n^{2})$ and space $O((U_{0}+U_{1})\cdot n)$, where $U_{0}=|\Sigma(\ell_{0})|$ and $U_{1}=|\Sigma(\ell_{1})|$ denote the number of distinct subset sums in each half of the input.
A randomized variant using the enumerator from Theorem 2 achieves expected runtime $O(\max(U_{0},U_{1})\cdot n)$ with the same space complexity.
Proof sketch.
The runtime follows by applying the analysis of Theorem 1 to each of the two halves of size $n/2$. The total work is governed by the larger of the two resulting enumerations due to the solver’s interleaved execution model, as justified in the subsequent section. ∎
5.2 Justification for the $\max(U_{0},U_{1})$ Runtime Bound in the Interleaved Model
Classical meet-in-the-middle solvers construct full enumeration tables for both halves of the input, yielding a runtime proportional to $U_{0}+U_{1}$. In contrast, our solver executes an interleaved enumeration and merge process, where elements of $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$ are matched on-the-fly. We now justify the tighter runtime bound of $O(\max(U_{0},U_{1})\cdot n^{2})$ for the deterministic variant and $O(\max(U_{0},U_{1})\cdot n)$ for the randomized one.
Naive Bound.
A sequential evaluation that first constructs both $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$ in full would incur a cost of $O((U_{0}+U_{1})\cdot n^{2})$ in the deterministic model.
Stronger Claim.
Our solver avoids full tabulation and instead performs feasibility checks incrementally. The merge logic operates online, emitting each candidate from one half and immediately testing it against the current frontier from the other. This yields a tighter complexity bound:
$T(n)=O(\max(U_{0},U_{1})\cdot n^{2})$
in the deterministic model, and
$T(n)=O(\max(U_{0},U_{1})\cdot n)$
in the randomized setting, where canonicality checks are $O(1)$ in expectation.
Formal Justification.
The total computational cost arises from:
  • enumerating all $U_{i}$ unique subset sums in side $\ell_{i}$, with $O(n)$ cost per bitmask update,
  • executing a constant number of hash table or lookup operations per candidate sum,
  • performing canonical deduplication when necessary.
Because the algorithm proceeds in rounds, extending one side’s frontier and probing the other immediately, the overall control flow is governed by the larger of $U_{0}$ or $U_{1}$. The smaller half’s contribution is subsumed within the dominant half’s traversal. Thus, no full duplication of effort occurs, and the interleaving guarantees that the additive bound $U_{0}+U_{1}$ is overly pessimistic.
Contrast with Horowitz–Sahni.
Classical meet-in-the-middle (e.g., Horowitz–Sahni [8]) uses an offline, two-phase architecture: first build both lists, then merge. This necessitates $O(U_{0}+U_{1})$ full enumeration regardless of structure. Our model, by contrast, halts exploration early when feasible and avoids redundant enumeration, exploiting the fact that $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$ need not be fully materialized to solve the instance.
Conclusion.
The interleaved solver’s runtime is structurally governed by the maximum work performed on either half. This aligns with both our theoretical analysis and empirical benchmarks, and further emphasizes the adaptivity of IC-SubsetSum to asymmetric structure.
5.3 Worst-Case Runtime via Controlled Aliasing
In addition to its adaptive performance, IC-SubsetSum guarantees a strict worst-case improvement over Horowitz–Sahni, even on collision-free instances. This is achieved via a deterministic Controlled Aliasing rule: a structural redundancy technique that reduces the enumeration space without sacrificing correctness. We describe an enhanced version of this technique that ensures full correctness for both the decision and construction variants of Subset–Sum.
Aliasing Rule.
Let $S$ be split into halves $\ell_{0}$ and $\ell_{1}$. In each half, select a pair of distinct elements—an alias pair—e.g., $(x_{0},x_{1})\in\ell_{0}$. During enumeration, we syntactically substitute the value of $x_{1}$ with that of $x_{0}$:
$\text{If }x_{1}\in P,\quad\text{compute }\sigma(P)\text{ using }x_{0}\text{ in place of }x_{1}.$
This substitution affects only the computed value; the subset’s bitmask identity is preserved. The result is that multiple distinct subsets collapse onto the same sum, reducing the enumeration state space. As proven in Appendix E, this yields a bound of
$O^{*}\left(\tfrac{3}{4}\cdot 2^{n/2}\right)=O^{*}(2^{n/2-\log_{2}(4/3)})\approx O^{*}(2^{n/2-0.415}).$
Canonical Aliasing via Parallel Memoization Lanes.
To preserve correctness under aliasing, we introduce a 4-lane memoization model. Each sum $\sigma$ is now associated with four parallel canonical entries, one for each aliasing state of the subset that generated it:
  1. (i)
    Neither $x_{0}$ nor $x_{1}$ included,
  2. (ii)
    $x_{0}$ included only,
  3. (iii)
    $x_{1}$ included only,
  4. (iv)
    Both $x_{0}$ and $x_{1}$ included.
These lanes are disjoint: a new prefix competes only within its own lane for lexicographic minimality. Thus, all semantically distinct witnesses are retained. The memoization structure becomes:
$\texttt{Memo}[\sigma]\in\left(\{\bot\}\cup\{0,1\}^{n}\right)^{4},$
i.e., a 4-tuple of bitmasks per reachable sum.
Compensatory Expansion.
During the final merge phase, we evaluate all valid interpretations of aliased sums. Let $s_{L}\in\Sigma(\ell_{0})$ and $s_{R}\in\Sigma(\ell_{1})$ be enumerated sums, and let $(x_{0},x_{1})$ and $(y_{0},y_{1})$ be the alias pairs in $\ell_{0}$ and $\ell_{1}$, respectively.
Define corrected sums:
$s^{\prime}_{L}$ $=s_{L}-x_{0}+x_{1}\quad\text{(if $x_{1}$ was present)}$)
$s^{\prime}_{R}$ $=s_{R}-y_{0}+y_{1}\quad\text{(if $y_{1}$ was present)}.$) .
We check all combinations:
$s_{L}+s_{R}=t,\quad s^{\prime}_{L}+s_{R}=t,\quad s_{L}+s^{\prime}_{R}=t,\quad s^{\prime}_{L}+s^{\prime}_{R}=t.$
This ensures correctness of the decision logic under value collision.
Witness Reconstruction.
Each of the four lanes retains a canonical prefix corresponding to its aliasing state, allowing for unambiguous witness recovery by combining the correct bitmasks from each half.
Remark 2.
While we focus on aliasing pairs for simplicity, the underlying 4-lane correctness mechanism is a scalable principle. The technique can be extended by using multiple aliases to the same element (e.g., $x_{1}\to x_{0}$, $x_{2}\to x_{0}$), which can yield a greater search space reduction. For example, a trio of aliased elements provides a speedup of $\epsilon=1$. Each additional alias increases the number of “lanes” and the complexity of the compensatory checks by a constant factor, but the asymptotic complexity remains unchanged. Analyzing the optimal trade-off is a promising direction for future work.
6 Refined View of Fine-Grained Hardness
Our algorithm challenges the standard landscape of fine-grained reductions. It shows that many hardness assumptions implicitly rely on a promise of low collision entropy—a structural assumption that classical reductions neither justify nor quantify. We make this dependence precise.
6.1 The Collision-Free Promise
We begin by formalising a restricted slice of Subset–Sum where classical hardness results are still meaningful.
Definition 3 (Collision-Resistant Subset–Sum).
For a parameter $\delta\in(0,1]$, define the class of instances with high entropy as
[…]
This condition ensures that any adaptive or certificate-sensitive algorithm cannot short-circuit the instance via early success or structural pruning. In particular, it guarantees that no-solution instances still require a near-exhaustive traversal of the exponentially large certificate space. Thus, hardness is preserved not just in satisfiable cases, but in verifying infeasibility—a setting where adaptive enumeration offers no shortcut unless entropy is low.
6.2 Refining SETH Lower Bounds
[…]
Theorem 5 (Revised SETH Lower Bound under Entropy Constraints).
[…]
Proof sketch.
If $\epsilon<\delta$, then $O^{*}(2^{\epsilon n})=o^{*}(2^{\delta n})\leq o^{*}(U)$, violating the known SETH lower bound for enumerating $\Sigma(S)$ in the collision-free regime. (See Appendix F for a formal justification of the connection between decision and enumeration hardness in this setting.)
6.3 Revisiting Fine-Grained Reductions
Many standard fine-grained reductions—for example, the classic reduction from CNF-SAT to Multiprocessor Scheduling [6]—produce Subset–Sum instances by mapping boolean literals to powers-of-two. This encoding guarantees that every subset sum is unique, i.e., $U=2^{n}$.
[…]
Take-Away for Fine-Grained Hardness.
Our work shows that fine-grained hardness results for Subset–Sum based on SETH are not truly unconditional. These lower bounds implicitly apply only to instances that are collision-free—i.e., where the number of distinct subset sums $U$ is nearly maximal ($U\approx 2^{n/2}$). Many known reductions, such as those that encode literals via powers-of-two, deliberately create this regime by avoiding arithmetic collisions.
Future reductions must explicitly certify high entropy, or restrict their claims to the collision-free regime. Otherwise, adaptive algorithms like IC-SubsetSum will evade the stated runtime barrier on all structured inputs.
6.4 Guidelines for Future Reductions
To maintain the validity of fine-grained lower bounds, future reductions should satisfy at least one of the following conditions:
  1. G1.
    Collision-resistance certification. Show that the generated instance $(S,t)$ satisfies $|\Sigma(S)|\geq 2^{\delta n}$ for some fixed $\delta>0$.
  2. G2.
    Entropy-aware hardness. Prove hardness not for all of Subset–Sum, but for a carefully defined structural subclass.
  3. G3.
    Non-enumerative goals. If the task is to return a single solution or check feasibility without generating $\Sigma(S)$, ensure that pruning-based methods (like IC-SubsetSum) provably fail on the instance family.
Discussion.
This perspective suggests a simple yet powerful conclusion: any worst-case lower bound for Subset–Sum—or for downstream problems like Knapsack or Scheduling—that fails to account for collision entropy is incomplete. Structural hardness must be made explicit. Otherwise, adaptive algorithms like IC-SubsetSum will continue to evade the theoretical limits by exploiting instance compressibility.
7 Experimental Validation
To validate our theoretical claims, we benchmark our enumerator on synthetic instances where we can precisely control the input structure. We use $n=48$ to allow for complete enumeration runs on a commodity desktop CPU. Our hypothesis is that introducing any form of additive redundancy will cause an exponential collapse in $U=|\Sigma(S)|$, with the runtime of IC-SubsetSum closely tracking this collapse.
We test this by manipulating the input set $S$ along three axes known to induce subset sum collisions: numeric density, duplicate elements, and additive progressions. For each experiment, we run the full enumeration without early termination to measure the total cost of generating the certificate $\Sigma(S)$. Further details and commentary on these experiments are available in Appendix D.
  • Numeric density. Forcing values into a smaller bit-length $w$ creates collisions via the pigeonhole principle.
  • Duplicate elements. Introducing identical elements creates trivial arithmetic dependencies.
  • Additive progressions. Injecting structured sequences creates non-trivial, correlated dependencies.
Table 1: Effect of structure on the size of the distinct sum set $U$. Each entry shows the ratio of observed unique sums to the worst-case maximum $2^{n/2}$ for one half of the set.
Structural Knob Setting $\frac{U}{2^{n/2}}$
Numeric density ($n{=}48$) $w=32\to 24\to 16$ 1.000.840.0271.00$\to$ 0.84$\to$ $0$ 0.84 → 0.027
Duplicates ($n{=}48$) 0240$\to$ 2$\to$ 40 → 2 → 4 duplicates 1.000.570.321.00$\to$ 0.57$\to$ 0.321.00 → 0.57 → 0.32
Additive progressions ($n{=}48$) 1 seq (len 4) / 2 seq (len 4) 0.690.690.69 / 0.470.470.47
Results.
As expected, we observed dramatic variation in the size of $\Sigma(S)$ under different input distributions. Uniform instances often had $U\approx 2^{n/2}$, while arithmetic progressions and structured inputs had $U$ values several orders of magnitude smaller. This confirms that $U$ is not a function of $n$ alone and supports its use as a certificate length parameter.
In all cases, the measured runtime of IC-SubsetSum scaled proportionally with the observed $U$ (for a fixed $n$), which is consistent with the proven $O(U\cdot n^{2})$ complexity bound.
8 Related Work
Subset Sum Algorithms.
The classical Horowitz–Sahni meet-in-the-middle algorithm [8] remains the fastest known worst-case solver for Subset Sum, with $O^{*}(2^{n/2})$ runtime. Despite extensive efforts, no unconditional improvement has been found in general [4, 16]. Our algorithm IC-SubsetSum departs from this tradition by analyzing runtime as a function of the number of distinct subset sums, $U$, rather than the input size $n$.
A recent randomized algorithm by Bringmann, Fischer, and Nakos [4] achieves the first unconditional improvement over Bellman’s dynamic programming algorithm for Subset Sum, solving the target-constrained decision problem in $\widetilde{O}(|S(X,t)|\cdot\sqrt{n})$ time. Their approach focuses on computing reachable sums up to a target $t$, using algebraic and combinatorial techniques. Our algorithm differs in scope: it deterministically enumerates the full set of unique subset sums $\Sigma(S)$, supports canonical constructive witnesses, and adapts its runtime to the instance’s internal structure. While the two approaches solve different variants of Subset Sum, they are complementary in goal and methodology.
Instance Complexity and Adaptive Algorithms.
Instance Complexity was introduced by Orponen, Ker-i, Uwe and Watanabe [10] to capture the inherent difficulty of individual instances. Although the model is powerful in theory, it was widely regarded as non-constructive and unimplementable—a perspective detailed in surveys on the topic (e.g., [5])—offering no path to practical algorithms. This paper directly addresses that long-standing gap. Our result shows, for the first time, how to constructively generate IC certificates online with a runtime that scales with certificate size.
Fine-Grained Complexity.
The fine-grained lower bound framework [15] has produced tight runtime barriers for many problems under assumptions like SETH. Our results add a structural dimension to this picture, showing that hardness results for Subset Sum only apply to collision-free instances. The requirement that reductions certify low entropy has not been previously made explicit. This parallels ideas in compressibility-aware reductions [1], but our formalization via $\Sigma(S)$ and entropy is novel.
Collision Structure in Combinatorics.
The entropy of subset-sum collisions has been studied both in additive combinatorics and in cryptanalytic settings, where low-entropy instances enable significantly faster algorithms. Austrin et al. [2] analyze the structure of subset sums under anti-concentration assumptions, while Howgrave-Graham and Joux [9] apply collision-based meet-in-the-middle techniques to exploit redundancy in integer knapsacks, achieving runtimes below the classical $2^{n/2}$ bound in the randomized setting. These results reinforce the idea that the subset-sum solution space can often be compressed substantially, and that this compressibility—quantified via collision entropy—directly governs algorithmic complexity.
Similarly, classic results in additive combinatorics—such as those related to the Littlewood–Offord problem [14]—bound the number of subset-sums that can collide based on anti-concentration phenomena. Our approach is instance-level and algorithmically constructive: we build data structures that prune repeated sums in real time and link the entropy collapse directly to runtime savings.
9 Future Directions
Closing Algorithmic Gaps for Subset Sum.
Our results open two natural algorithmic questions. First, for our randomized $O(U\cdot n)$ algorithm, can the linear factor in $n$ be eliminated to achieve an $\tilde{O}(U)$ runtime? Second, and more fundamentally, can our deterministic $O(U\cdot n^{2})$ algorithm be improved to match the randomized bound, perhaps by finding a deterministic method for canonical comparison that avoids the $O(n)$ cost?
Structure-Sensitive Runtimes in NP.
Can other NP-complete problems admit certificate-sensitive runtimes? This paper introduces a new structural knob—collision entropy—that complements traditional parameters like solution density. For Partition, Knapsack, and Bin Packing, it may be possible to construct a compressed summary of feasible configurations in time proportional to its size. Formalising such certificates and bounding the corresponding enumeration complexity are exciting open problems.
Derandomization of Local Canonicality.
The randomized $O(U\cdot n)$ algorithm relies on pairwise polynomial hashing to simulate lexicographic comparisons in expected constant time. Is there a deterministic analog that achieves sublinear-time prefix comparison in the local model? Any such result would not only match the conjectured optimal runtime but also resolve a form of local derandomization in a highly practical setting.
Constructive Instance Complexity Beyond Subset Sum.
This paper builds a bridge between Orponen-style instance complexity and efficient enumeration. For Subset–Sum, we show that the certificate $\Sigma(S)$ can be constructed in deterministic $O(U\cdot n^{2})$ time or expected $O(U\cdot n)$ time. Can similar certifiers be constructed online for problems like 3SAT or Clique, and if so, under what entropy or compressibility assumptions? A broader theory of algorithmic IC—one that tracks both information content and construction cost—remains to be written.
Toward a Complexity Class for Certificate-Sensitive Problems.
Our algorithm solves Subset Sum in time polynomial in both the size of its minimal instance-specific certificate, $U=|\Sigma(S)|$, and the input size $n$. This motivates a broader question: are there other NP-complete problems that admit solvers whose runtime is polynomial in the size of an intrinsic structural certificate of the instance? We tentatively propose the name P-IC (Polynomial in Instance Certificate) for such problems. Formalizing this class would require a general definition of “instance certificate”—for example, the set of all satisfying assignments in SAT, or the set of all achievable (value, weight) pairs in Knapsack. Unlike parameterized complexity (e.g., FPT), where the parameter is typically a small, fixed integer supplied with the input, the certificate in our setting is an emergent property of the instance’s internal structure, not known a priori. We believe that defining the boundaries of P-IC, and identifying other problems that fall within it, is a promising direction for future work.
10 Conclusion
Theoretical Advance.
This paper develops a new framework for structure-aware enumeration in NP-complete problems, with Subset Sum as the case study. We introduce a certificate-sensitive complexity measure $U=|\Sigma(S)|$ and show that it aligns with both instance difficulty and practical runtime. Our results offer a unified explanation for why real-world instances are easier, why SETH-based reductions can be structurally brittle, and how instance complexity can be made constructive.
Algorithmic Advance.
IC-SubsetSum achieves a set of performance guarantees that collectively redefine the state of the art:
  • Certificate-sensitive runtime. A deterministic worst-case bound of $O(U\cdot n^{2})$ and an expected randomized bound of $O(U\cdot n)$, where $U$ is discovered online.
  • Worst-case improvement over classic methods, achieving $O^{*}(2^{n/2-\varepsilon})$ for some fixed $\varepsilon>0$.
  • Real-time certificate generation, producing the minimal witness $\Sigma(S)$ with zero redundancy.
Closing the gap between the deterministic and randomized bounds remains a key open challenge, as does the question of whether an $\tilde{O}(U)$ runtime is achievable.
May all your programs be short and quick (Fortnow [5]) and may their runtimes reveal the structure within.
Appendix Overview
Provenance.
This appendix material expands on the enumeration logic, correctness conditions, and structural optimizations introduced across Sections 2–4 and Section 6 of the main paper [11].
We provide the following appendices:
Appendix Overview
This appendix material expands on the enumeration logic, correctness conditions, and structural optimizations introduced in Sections 46 of the main paper.
  • Appendix A provides a detailed description and pseudocode for the core column-wise subset sum enumerator.
  • Appendix B lists the algebraic combination lemmas used by the solver, including correctness checks for sum recombination.
  • Appendix C describes the full double-meet-in-the-middle framework used to exploit structure and symmetry.
  • Appendix D presents additional experimental data, including plots of runtime scaling and collision entropy.
  • Appendix E presents the full Controlled Aliasing framework, including the 4-lane memoization model and implementation-level pseudocode.
  • Appendix F provides a formal justification for the connection between decision and enumeration hardness in high-entropy instances.
Appendix A Column-wise Enumerator: Full Details
Algorithm Description
Column-wise expansion.
Our state space consists of all canonical partial sums $\sigma$ that are still under construction. At each step, we extend $\sigma$ by adding one unused element $a_{i}\notin R$, and we update the prefix to $R\cup\{a_{i}\}$ if the resulting sum $\sigma^{\prime}=\sigma+a_{i}$ is new or establishes a more canonical representation.
This leads to at most $O(U\cdot n)$ total extension attempts from canonical states. The final complexity, as established in Section 4.1, depends on the cost of the canonical comparison performed during collisions.
High-density handling.
To guarantee correctness in high-collision regimes, we use a canonicalization rule based on the lexicographic order of index bitmasks. This ensures that each sum $\sigma$ is associated with exactly one canonical prefix, preventing incorrect pruning due to non-deterministic resolution of duplicated sums.
4-Column Litmus Test.
The real-time nature of our collision detection enables a simple and effective litmus test to estimate an instance’s hardness. By running the enumerator for only a few initial columns (e.g., up to $k=4$)—a process that runs in polynomial time $O(n^{4})$—we can measure the early-stage collision rate among small subsets. A high collision rate strongly predicts that the total number of unique subset sums $U$ will be small, signaling a structurally “easy” instance. Conversely, a low collision rate suggests that $U$ will be large, indicating a harder, unstructured input. This test offers a fast, lightweight method to assess an instance’s intrinsic complexity before engaging in a full exponential-time computation.
A.1 Pseudocode
Algorithm 1 EnumerateUniqueSubsetSums
1:Set $S=\{a_{1},\dots,a_{n}\}$
2:Memo table mapping each unique subset sum to its canonical bitmask.
[…]
5:for $k=1$ to $n/2$ do
[…]
8:   for $i=1$ to $n$ do
9:     if bit $i$ is not set in mask then
10:      new_sum $\leftarrow\texttt{sum}+a_{i}$
11:      new_mask $\leftarrow\texttt{mask}\lor(1\ll(i-1))$
12:      if new_sum $\notin$ Memo then
13:        Memo[new_sum] $\leftarrow$ new_mask
[…]
15:      else if IsLexicographicallySmaller(new_mask, Memo[new_sum]then
16:        Memo[new_sum] $\leftarrow$ new_mask
[…]
18:  Frontier $\leftarrow$ NextFrontier
19:  if Frontier =$\emptyset$= ∅ then break
20:return Memo
Algorithm 2 IsLexicographicallySmaller
[…]
2:  for $i=1$ to $n$ do
3:   $a_{i}\leftarrow$ bit $i$ of $A$
4:   $b_{i}\leftarrow$ bit $i$ of $B$
5:   if $a_{i}<b_{i}$ then return true
6:   else if $a_{i}>b_{i}$ then return false
7:  return false $\triangleright$ They are equal
A.2 Implementation Invariants
To guarantee correctness and optimal performance, the enumerator must uphold:
  • Deterministic tie-breaking. Among colliding prefixes for a given sum $\sigma$, the lex-minimal one is always retained. This guarantees canonicity across all layers.
  • Efficient storage. Total space is dominated by storing the $U$ canonical prefixes, each represented by an $n$-bitmask. This leads to a space complexity of $O(U\cdot n)$.
  • Duplication elimination. Every sum $\sigma\in\Sigma(S)$ is generated exactly once. This avoids wasted work in high-density regimes and maintains instance-optimality.
Appendix B Compositional Sum Lemmas and Complement Check
These lemmas govern how partial sums are combined across recursive calls and meet-in-the-middle partitions. They apply to both aliasing and non-aliasing regimes and are invoked by all canonical certificate routines to ensure uniqueness and correctness.
Let $\ell_{0},\ell_{1}$ be the two halves of a split of $S$, and write $\Sigma(\ell_{i})$ for the set of distinct subset sums produced by the enumerator on side $i$.
Lemma 1 (Direct Match in One Half).
[…]
$\sigma(A)=t.$
Lemma 2 (Self-Complement Match).
[…]
$\sigma(\ell_{i}\setminus A)=\sigma(\ell_{i})-t.$
Lemma 3 (Cross-Half Direct Match).
For subsets $A\subseteq\ell_{0}$ and $B\subseteq\ell_{1}$,
$\sigma(A)+\sigma(B)=t\quad\Longleftrightarrow\quad t-\sigma(A)\in\Sigma(\ell_{1}).$
Lemma 4 (Double Complement Match).
If subsets $A\subseteq\ell_{0}$ and $B\subseteq\ell_{1}$ satisfy
$\sigma(\ell_{0}\setminus A)+\sigma(\ell_{1}\setminus B)=\sigma(S)-t,$
then $S\setminus(A\cup B)$ is a witness.
Lemma 5 (Mixed Case Match).
A subset $A\subseteq\ell_{0}$ and the complement of a subset $B\subseteq\ell_{1}$ form a solution if and only if
$\sigma(A)+\sigma(\ell_{1}\setminus B)=t.$
Proof sketch.
All statements follow from elementary algebra on disjoint subsets and set complements. These identities justify the conditional checks performed by the Check routine in the main solver. Since the enumerator generates all elements of $\Sigma(\ell_{i})$ correctly (§A), all valid witnesses are discoverable by the algorithm.
Full derivations appear in §6.1–6.3 of [11]. ∎
Appendix C Double-MIM Solver: Detailed Pseudocode and Analysis
C.1 Algorithm Sketch
We assume the set $S=\{a_{1},\dots,a_{n}\}$ is split into left and right halves, $\ell_{0}$ and $\ell_{1}$. While this partition can be arbitrary, its balance can be optimized beforehand via a lightweight, polynomial-time analysis of the instance’s local collision structure (4-Column Litmus Test, Appendix A). We apply the unique-subset-sum enumerator (Appendix A) to $\ell_{0}$ and $\ell_{1}$ to construct their respective certificates, $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$. Then, instead of using a traditional sorted merge, solutions are checked in real time: as each new unique sum is discovered, the algorithm applies the Combination Lemmas and Complement Logic (Appendix B) to determine whether the new sum, or its complement, combines with any previously seen sum from the other split to reach the target $t$.
Algorithm 3 Certificate-Sensitive Subset Sum Solver
1:Multiset $S=\{s_{1},\dots,s_{n}\}$, target $t$
2:Returns a certificate if $t\in\Sigma(S)$; otherwise reports failure
3:Split $S$ into $\ell_{0},\ell_{1}$ alternating by index
4:Input $\leftarrow\{\mathtt{kPerm}(0,\mathbf{0},b)\mid b\in\{0,1\}\}$
[…]
7:for $r=0$ to $n/4$ do
[…]
9:  for all $p\in$ Input do
10:   $L\leftarrow\ell_{p.\texttt{split}}$
[…]
14:     if $p.\texttt{elements}[i]=0$ then
15:      $s^{\prime}\leftarrow p.\texttt{sum}+L[i]$
16:      $e^{\prime}\leftarrow p.\texttt{elements}$ with bit $i$ set to 111
17:      $q\leftarrow\mathtt{kPerm}(s^{\prime},e^{\prime},p.\texttt{split})$
18:      if $s^{\prime}\notin$ Memo then
[…]
21:        NextFrontier $\leftarrow$ NextFrontier $\cup\{q\}$
[…]
24:        Remove $\texttt{Cert}[s^{\prime}]$ from NextFrontier
[…]
27:        NextFrontier $\leftarrow$ NextFrontier $\cup\{q\}$
28:  Input $\leftarrow$ NextFrontier
29:return ”No solution found”
Real-Time Solution Check Procedure (invoked during enumeration):
[…]
2:Verifies whether $q$ completes a valid solution to the target
3:Let $t$ be the target
[…]
7:if $s^{\prime}=t$ or $Sum_{b}-s^{\prime}=t$ then
8:  Output: direct solution found
9:  return
10:if $t-s^{\prime}\in$ Memo then
11:  Output: solution across both splits
12:  return
13:if $t-(Sum_{b}-s^{\prime})\in$ Memo then
14:  Output: solution via complement merge
15:  return
C.2 Time and Space Bounds
Let $U_{0}$ and $U_{1}$ be the number of distinct subset sums in the two halves, $\ell_{0}$ and $\ell_{1}$.
Time Complexity.
The runtime of the full solver is governed by the unique-subset-sum enumerator applied to each half of the input. As established in Theorem 1, the deterministic enumerator performs $O(U_{\text{local}}\cdot k^{2})$ work on a set of size $k$ with $U_{\text{local}}$ unique sums. Since our splits are of size $n/2$, each half incurs $O(U_{i}\cdot(n/2)^{2})$ work in the worst case. Although both sets of unique subset sums, $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$, are retained in memory to support canonical certificate construction, the runtime is asymptotically dominated by the larger enumerator. This yields a total deterministic complexity of
$O(\max(U_{0},U_{1})\cdot n^{2}).$
The randomized algorithm achieves expected runtime
$O(\max(U_{0},U_{1})\cdot n),$
as shown in Theorem 2.
Space Complexity.
The space is dominated by storing the canonical prefixes for each of the $U_{0}$ and $U_{1}$ states. Each prefix is an $n/2$-bitmask. Therefore, the total space complexity is $O((U_{0}+U_{1})\cdot n)$.
Optimality.
Our randomized algorithm has expected runtime $O(\max(U_{0},U_{1})\cdot n)$, matching the conjectured $\Omega(U\cdot n)$ lower bound from Conjecture 1. This establishes its conditional optimality in expectation. The deterministic algorithm achieves $O(\max(U_{0},U_{1})\cdot n^{2})$ runtime, leaving a polynomial gap relative to the conjectured lower bound. Closing this gap with a deterministic certificate-sensitive algorithm remains an open problem.
Appendix D Experimental Details
We provide additional data and commentary on the experiments presented in Section 7. Our focus is on how structure (e.g., low entropy or repeated values) affects the size of the certificate $\Sigma(S)$, and how this correlates with observed runtime.
D.1 Extremely dense instances ($n=100$)
[…]
The data clearly show that the number of distinct subset sums $U$ falls far below the $2^{n/2}$ barrier as $w$ decreases, and the runtime is significantly reduced once $U\ll 2^{n/2}$, consistent with the proven $O(U\cdot n^{2})$ certificate-sensitive bound.
Full experimental results, implementation details, and benchmark data can be found in Section §6 of [11].
D.2 Summary
Across all experiments, we observe that the certificate size $U$ varies by orders of magnitude depending on the input distribution, and the enumerator’s wall-time closely tracks the drop in $U$. This validates the $O(U\cdot n^{2})$ runtime scaling proven in Section 1, as performance for a fixed $n$ is proportional to $U$.
The experiments confirm that IC-SubsetSum is highly adaptive in practice, with performance governed not by the input size alone, but by the structural complexity of the instance itself.
Appendix E Controlled Aliasing: Correctness and Implementation
To ensure correctness under value aliasing, we introduce a 4-lane memoization scheme. Each reachable sum stores up to four canonical bitmasks, corresponding to distinct aliasing configurations. This enables both decision and construction variants of Subset–Sum to be solved without error, even when synthetic collisions are introduced.
Memoization Structure
We define the memoization table as:
$\texttt{Memo}:\mathbb{Z}\rightarrow\left(\{\bot\}\cup\{0,1\}^{n}\right)^{4}$
Each entry maps a sum $\sigma$ to an array of four bitmasks:
[…]
State Space Reduction via Aliasing
Lemma 6 (State Space Reduction via Aliasing).
Let $S^{\prime}\subseteq\mathbb{Z}_{>0}$ be a set of $k$ positive integers, and let $x_{0},x_{1}\in S^{\prime}$ be a designated aliasing pair with $x_{0}\neq x_{1}$. Suppose we apply the aliasing rule during subset sum enumeration: whenever $x_{1}$ appears in a subset, its value is treated as $x_{0}$. Then the total number of distinct subset sums generated is at most $\frac{3}{4}\cdot 2^{k}$.
Proof.
Partition the $2^{k}$ subsets of $S^{\prime}$ according to whether they include neither, one, or both of $x_{0}$ and $x_{1}$. Let $P\subseteq S^{\prime}\setminus\{x_{0},x_{1}\}$ be an arbitrary subset of the remaining elements.
There are four corresponding supersets of $P$ formed by optionally adding $x_{0}$ and/or $x_{1}$:
  • $P$
  • $P\cup\{x_{0}\}$
  • $P\cup\{x_{1}\}$
  • $P\cup\{x_{0},x_{1}\}$
Under normal enumeration, each contributes a distinct sum. Under the aliasing rule, $x_{1}$ is treated as $x_{0}$, so the last three all contribute sums of the form $\sigma(P)+x_{0}$, except the last, which contributes $\sigma(P)+2x_{0}$. Thus, the four subsets yield only three unique sums. Since there are $2^{k-2}$ such base subsets $P$, the total number of unique subset sums is at most $3\cdot 2^{k-2}=\frac{3}{4}\cdot 2^{k}$. ∎
Core Procedures
Alias State Function.
[…]
2:  return $2\cdot b[i_{1}]+b[i_{0}]$
Insert into Memo Table.
[…]
2:  $a\leftarrow\textsc{AliasState}(b,i_{0},i_{1})$
[…]
Corrected Sum.
[…]
2:  if $a=2$ or $a=3$ then
3:   return $\sigma-x_{0}+x_{1}$
4:  else
5:   return $\sigma$
Aliased Value (Inverse).
[…]
2:  if $a=2$ or $a=3$ then
3:   return $\sigma_{real}+x_{0}-x_{1}$
4:  else
5:   return $\sigma_{real}$
Compensatory Witness Recovery
To ensure correctness during merging, the algorithm performs a 16-way compensatory check across alias states of both halves. For each left-side sum, it computes the required right-side partner under all alias state interpretations.
FindSolution Routine.
[…]
2:  for all $\sigma_{L}\in\texttt{MemoLeft}$ do
3:   for $a_{L}=0$ to 333 do
4:     $b_{L}\leftarrow\texttt{MemoLeft}[\sigma_{L}][a_{L}]$
5:     if $b_{L}=\bot$ then continue
6:     $s_{L}\leftarrow\textsc{CorrectedSum}(\sigma_{L},a_{L},x_{0},x_{1})$
7:     $s_{R}\leftarrow t-s_{L}$
8:     for $a_{R}=0$ to 333 do
9:      $\sigma_{R}\leftarrow\textsc{AliasedValue}(s_{R},a_{R},y_{0},y_{1})$
10:      if $\sigma_{R}\in\texttt{MemoRight}$ then
11:        $b_{R}\leftarrow\texttt{MemoRight}[\sigma_{R}][a_{R}]$
12:        if $b_{R}\neq\bot$ then
13:         return $b_{L}\,\|\,b_{R}$
14:  return None
Complexity Summary
  • Each reachable sum stores up to 4 bitmasks.
  • All routines operate in constant time per alias lane.
  • Total time and space remain:
[…]
  • Merge routine performs $O(U)$ lookups with constant overhead.
Conclusion
This 4-lane structure solves the canonical aliasing problem completely. It preserves semantic distinctions between aliased subsets, maintains lexicographic minimality in each lane, and supports full witness reconstruction without loss of correctness or efficiency. The Controlled Aliasing mechanism now guarantees correctness for both decision and construction variants of Subset–Sum on all inputs.
Appendix F Note: Decision vs. Enumeration Hardness
[…]
  • Perform a prefix-guided traversal of the subset tree using $\mathcal{A}$ as an oracle.
  • At each node, extend the prefix by one bit (include or exclude the next element).
  • Query whether any extension leads to a feasible sum in the instance.
  • Prune infeasible branches, recurse on feasible ones, and memoize discovered sums.
[…]